OpenStudy (anonymous):

2H2O(l)--> 2H2(g)+O2(g) If 50 grams of liquid water were completely changed to hydrogen and oxygen gas, how many grams of gas should there be produced? A) 25 B) 75 C) 50 D) 100 (I know it is not 100 or 25)

OpenStudy (aaronq):

you need to do some stoichiometry. In stoichiometry, the calculations are always* done with moles, so the first step is always: 1. Convert what you have to moles (*sometimes you can use volumes when working with gases) For your example, you use: \(\sf moles=\dfrac{mass}{Molar~ mass}\) 2. Set up a ratio to relate the moles of reactant to moles of product using the stoichiometric coefficients (i.e. molar ratio). \(\sf \dfrac{moles~of~H_2O}{H_2O's~coefficient} =\dfrac{moles~of~H_2}{H_2's~coefficient} \) plug in what you know then solve. NOTE: you need to do this for both \(H_2\) and \(O_2\) since they're both gases. 3. Convert moles to mass using the formula from 1. For this exercise, you need to add the masses of both gases.

OpenStudy (anonymous):

Thank you!! @aaronq